A true-false test has ten questions. A student guesses all the answers. Find the probability of getting at least eight questions correct.
Answer by stanbon(75887)

A true-false test has ten questions. A student guesses all the answers. Find the probability of getting at least eight questions correct.
------------------------------
Binomial Problem:
n = 10 ; p = 1/2 ; 8<= x <=10
--------------------------------
P(8<= x <=10) = 1 - binomcdf(10,1/2,7) = 0.0547
===================================================
Cheers,
Stan H.